In 1991, the life expectancy of males in a certain country was 66.1 years. In 1998, it was 68.5 years. Let E represent the life expectancy in year t and let t represent the number of years since 1991. the linear functin E (t) that fits the dat is:
E (t) = __t +___.
(round to the nearest tenth)
Use the function to predict the life expectanc of males in 2008.
E (17) = ____
(round to the nearest tenth)

Answer by stanbon(75887)

In 1991, the life expectancy of males in a certain country was 66.1 years. In 1998, it was 68.5 years. Let E represent the life expectancy in year t and let t represent the number of years since 1991.
---
2 points: (0,66.1), (7,68.5)
---
slope = (68.5-66.1)/7 = 12/35
int: 66.1
------------------------------------
the linear function E (t) that fits the date is:
E (t) = (12/35)t +66.1
(round to the nearest tenth)
----------------------------------------------
Use the function to predict the life expectanc of males in 2008.
E (17) = (12/35)17+66.1 = 71.93
(round to the nearest tenth)
===================================
Cheers,
Stan H.
=======